how much heat is needed to melt 1.5kg of ice and then to rise the temperature of the resulting water 500c?
1
Expert's answer
2020-06-08T10:25:39-0400

The amount of heat required to melt ice and raise the temperature of water "T^oC" is

"Q=mL_f+mc\Delta T"

Here "m=1.5 kg"

"L_f=3.33*10^5 J\/kg" is Latent heat of fusion of ice.

"c=4186 J\/kg.^oC" is heat capacity of water.

"\Delta T=50 ^oC-0^oC=50^oC"

So,

"Q=(1.50)[3.33*10^5 +(4186)(50)]=8.13*10^5 J"